OpenStudy (anonymous):

The question is "What is one of the vertices of the hyperbola (y+1)^2 - (x+1)^2/4 =1 Please help me and thanks in advance :)

OpenStudy (anonymous):

do you know where the center of this hyperbola is?

OpenStudy (shubhamsrg):

you're welcome in advance too! ;)

OpenStudy (anonymous):

Umm no?? I bearly started taking this class today like for 2 mins :(

OpenStudy (anonymous):

How would you find the center?

OpenStudy (anonymous):

what does your notes/textbook say about the standard equation of a hyperbola?

OpenStudy (anonymous):

the equation tells you the center...

OpenStudy (anonymous):

x^2/a^2 -/+ y^2/b^2=1

OpenStudy (anonymous):

we need the more general form.. lemme put it up...

OpenStudy (anonymous):

\[\large \frac {(x-h)^2}{a^2}- \frac {(y-k)^2}{b^2}=1\] \[\large \frac {(y-k)^2}{a^2}- \frac {(x-h)^2}{b^2}=1\]

OpenStudy (anonymous):

in either case, the center is (h,k). now can you identify your center?

OpenStudy (anonymous):

It would be (1,1) right?

OpenStudy (anonymous):

not quite.. it's (-1, -1)

OpenStudy (anonymous):

Is that because they were being subtracted?

OpenStudy (anonymous):

in the formula, yes... and since your equation is (x+1)^2 you can write it as (x - (-1))^2 understand?

OpenStudy (anonymous):

Yes. Thanks make so much more sense now :)

OpenStudy (anonymous):

once you have the center, we can find your vertices to answer your question.

OpenStudy (anonymous):

So thats when we plug in our vertices into the equation?

OpenStudy (anonymous):

well, we need to find the vertices...

OpenStudy (anonymous):

the equation actually tells us WHERE the vertices are...

OpenStudy (anonymous):

in relation to the center..

OpenStudy (anonymous):

Oh I got you :)

OpenStudy (anonymous):

Thanks DPALNC":)

OpenStudy (anonymous):

first... i wrote down two types of equations... do you know why?

OpenStudy (anonymous):

yw...

OpenStudy (anonymous):

depends if its vertical or horizantal

OpenStudy (anonymous):

horizontal*

OpenStudy (anonymous):

correct... what type do you have?

OpenStudy (anonymous):

Vertical

OpenStudy (anonymous):

correct...

OpenStudy (anonymous):

so all you need to know is what is your denominator under the (y+1)^2 term... that is a^2. what is a^2 and what is a ?

OpenStudy (anonymous):

Would it be 4?

OpenStudy (anonymous):

that's under the (x+1)^2... what's under (y+1)^2 ?

OpenStudy (anonymous):

There is nothing under it.... Would there be like an invisible one?

OpenStudy (anonymous):

Or what would take place there?

OpenStudy (anonymous):

yes.. invisible 1

OpenStudy (anonymous):

so a=1

OpenStudy (anonymous):

So now it would be 1,-1

OpenStudy (anonymous):

because you said we have the vertical branch hyperbola, from the center of (-1, -1), you go UP 1 and DOWN 1 from the center. can you now give me the coordinates of the vertices?

OpenStudy (anonymous):

 there's your two vertices...